Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
fiori, tringali, berni, gherardini, nardecchia, spinicelli - 18:59 Saturday 09 March 2024 (63540) Print this report
Environmental shift - magnetic injection map, 25-min glitch investigation, new mysterious noise at NEB

Magnetic injection map:

We performed a new set of near field magnetic injection to map the local couplings of the NI, WI, BS, SIB1, SDB1 towers

The plan was to do injections and terminals but we could not because the ITF repetedly unlocked due to a (new) mysterious noise occurring at the NEB

The attached file gives gps times and details.

25-min glitch investigation:

We perfomed the following actions

  • unplugged the power supplier and pre-amplifier of the NI point absorber (TCS room)
  • unplugged thepoint absorber pico-motors driver
  • unplugged the thorlab module tri-axis piezo controller (piezo driver laser CO2)
  • physically disconnected all the cables from the point absorber module (NI tower base)
  • we left opended for the full shift duration (> 8 hours ) the door of the TCS rooom. Figure 1 shows the evolution of the temeprature of the TCS room and the PR CHRoCC parameter

Glitches are still there :-((

Any effect on the glitch rate has to be better checked

A new mysterious transient noise at NEB:

As described by Fabio in  63536, today starting from around 12:00 UTC the ITF unlocked several times because the NE suspension F0 got excited. A correlated noise is present in several ENV and NN sensors in NEB: accelerometers and microphones

From a preliminary analysis we notice that

  • the noise is transient with a duration of a few seconds, the intensity of the noise is not always the same
  • the noise has a peculiarr shape in the spectrogram, consisting of a peak excitation at few hundred of Hz followed by a "whisker": see Figure 2 and 3
  • the noise source seems located inside the NEB experimental room, because looking at the NN microphonic array the mics placed inside the external technical rooms (number 6, 7 and 8) do not see the noise: Figure 4
  • looking at the NN seismic array, the noise is not uniformely distributed inside  the NEB hall, but is seems less intense at the far end of the room (towards the tunnel) and more intense along the west side of the room: Figure 5
  • The rate of these glitches is not regular and not always caused an unlock: Figure 6 and Figure 7 (zoom)
  • we also checked the NEB environmental dataset of the whole day, and it loooks that the noise first appeared around 11:45 UTC
Images attached to this report
Non-image files attached to this report
Comments to this report:
Tringali, Gherardini - 12:47 Sunday 10 March 2024 (63545) Print this report

The audio files of this new noise are generated using the microphone located in the NE experimental hall, ENV_NEB_MIC, during the following times:

Mar 9, 12:04:20 UTC, gps=1394021078 , dur=50 sec;
Mar 9, 12:15:20 UTC, gps=1394021738 , dur=50 sec;
Mar 9, 15:53:10 UTC, gps= 1394034808, dur=50 sec;
Mar 10, 07:17:20 UTC, gps=1394090258, dur=50sec.

Non-image files attached to this comment
Gherardini, Vacuum - 18:44 Sunday 10 March 2024 (63547) Print this report

Today at around 4 pm, we turned off the turbo of the NE tower, suspected of producing the recorded noises (possible controller malfunction)

berni - 22:47 Sunday 10 March 2024 (63549) Print this report

To be checked the unlock from CARM_NULL_1F (shutter open) occurred at 18:50 UTC

Images attached to this comment
fiori - 23:24 Sunday 10 March 2024 (63550) Print this report

From a quick look, considering the excited frequencies and the concerned sensors, I would say it is very similar to the "mysterious" noise.

Figure 1 shows the event of 18:50 UTC (pointend out by Francesco)

Figure 2 shows the previous event (15:05 UTC) (double) which was followed right after by the action by Antonio.

Images attached to this comment
Paoletti - 10:19 Monday 11 March 2024 (63553) Print this report
Images attached to this comment
Paoletti - 14:10 Tuesday 12 March 2024 (63568) Print this report

Answering a question that arose during today's commissioning meeting: the TCS room temperature variation (1.6 degC) obtained opening the door was performed while the ITF was having lock problems (NEB noise), so there are few glitch samples available for refined analysis.

However, looking at past data, we can see that the TCS room already had the same variations, so we can analyze the old data to find a correlation with the glitch rate at 25m.

Images attached to this comment
Vacuum - 19:23 Thursday 14 March 2024 (63614) Print this report

as a followup of the investigations of the new noise at NE,  during last tuesday break - 12/03 morning - we did the following :

- restarted the main turbopump (400 Hz)
- stopped  the 'spare' pump (820 Hz) which has been in operation for months

Gherardini, Vacuum - 22:38 Saturday 16 March 2024 (63645) Print this report

warned by Fabio, at around 1915LT  we again switched OFF the main turbopump at NE and switched ON the spare one. Apparently the pump did not emit any alarm, anyway the sound recordings are likely from a malfunctioning turbo pump.

In practice, this is the same configuration tested in 63550 .

Note: the spare pump rotating frequency has been set at 547 Hz

fiori, tringali, pasqualetti, gherardini, magazzù - 15:55 Sunday 17 March 2024 (63647) Print this report

"NE mysterious noise"

One additional event occurred March 16 at 22:23 UTC and caused an unlock: Figure 1 (trend), Figure 2 and Figure 3 (spectrogram of accelerometers). The frequency of the turbo pump line (547Hz) did not seem to be affected at all (Figure 4). 

At that time, according to Antonio's and Fabio's reports: the main turbo pump was switched off and the spare turbo pump was on, the accelerometer named ENV_NE_BOTTOM_ACC was placed on the main turbo (see picture in Figure 5).

Figure 6 compares the spectrum of this accelerometer (orange) with similar accelerometers on the NE big flange (blue) and on the NE cryo trap (green). The signal seen by the accelerometer on the main pump indeed  detected a large signal.

We moved the accelerometer named ENV_NE_MIDDLE_ACC on top of the spare turbo pump. See picture in Figure 7.

We switched off the external cooling fan of the main turbo.

Extra glitches in accelerometer ENV_NE_ACC_Z:  in Figure 2, we notice several fast glitches which are not seen by the other accelerometer. Need further investigation. 

Images attached to this comment
salvador - 11:53 Monday 18 March 2024 (63655) Print this report

I've added the "INF_TCSroom_TE" channel to the previously shown data of the glitch period. Fig. 1 shows the most correlated temperature channels with the glitch period. The channels have been time shifted according to what Francesco extracted (
The shif of this new channel is around 7 h before the glitch period making it even earlier than "INF_NI_BOTTOM_TE1" by almost 3 h. Although, the correlation seems smaller compared to the "BOTTOM_TE1" channel.

Fig. 2 shows a close-up of these channels.

Images attached to this comment
Vacuum - 18:24 Monday 18 March 2024 (63666) Print this report

This afternoon we switched off the controller of the main turbo at NE, suspected of somehow producing disturbances even when the pump is stopped .

Search Help